2. A company with a 10.5% required return is considering a new project which will affect its net working capital requirements over the life of the project. The company’s marginal tax rate is 21%. What is the effect on NPV of the net working capital requirement?

0 1 2 3 4 5
Change in NWC 45,000 6,750 7,763 8,926 10,266 -78,705

Now assume that the company implements a lean manufacturing initiative that cuts by half the numbers in the above table. How would this affect the NPV of the project? Is this a good idea?
Answer #1

NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= -45000-6750/1.105^1-7763/1.105^2-8926/1.105^3-10266/1.105^4 +78705/1.105^5

=-23193.82

At 50% of the above

NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= -45000/2-(6750/2)/1.105^1-(7763/2)/1.105^2-(8926/2)/1.105^3-(10266/2)/1.105^4 +(78705/2)/1.105^5

=-11596.91

This reduces the decrease in NPV to half and hence is a good idea.